# Leveraging Soft Prompts for Privacy Attacks in Federated Prompt Tuning

### v1

@@ 3.2.2 PromptMIA
To build up on and overcome the limitations of the naive attack, we impose two requirements on adversarial key generation. First, adversarial keys must achieve higher cosine similarity with the target query $q(\mathcal{T})$ than any benign key. Second, adversarial keys must be sufficiently diverse from one another to not be trivially exposed. Formally, the server selects a subset $S\subseteq[M]$ of size $N$, and for each index $j\in S$, the original key $k_{j}$ is replaced with an adversarial key $k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}$ ( see Eq. 6). $k_{a_{m}}$ is constrained to lie within a cosine similarity interval with $q(\mathcal{T})$:
$\kappa\!\left(q(\mathcal{T}),k_{a_{m}}\right)\in\Bigl[\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)+\delta_{\min},\;\;\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)+\delta_{\min}+\Delta\Bigr],\forall\,k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}$ (7)
where $\delta_{\min}$ ensures that $\kappa\!\left(q(\mathcal{T}),k_{a_{m}}\right)>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)$, while $\Delta$ introduces controlled variability to prevent all adversarial keys from collapsing to the same vector. Algorithm 1 describes the basic procedure for generating a single adversarial key with a specified cosine similarity $s$ to the target query. Building on this primitive, Algorithm 2 constructs a set of $N$ adversarial keys $\mathcal{K}_{\textsc{ADV}}$ that satisfies Eq. 7. Details of these algorithms are given in Appendix A.2. PromptMIA allows the adversary to generate an adversarial key set $\mathcal{K}_{\textsc{adv}}$ that will always be selected and updated when $\mathcal{T}\in\mathcal{D}$, while not being easily detectable (see Fig. 2(c) and Section 4.2). The server generates and distributes the modified global prompt pool $(\tilde{\mathcal{K}},\tilde{\mathcal{P}})=\bigl(\mathcal{K}_{\textsc{adv}}\cup\mathcal{K}_{\textsc{benign}},\;\mathcal{P}_{\textsc{adv}}\cup\mathcal{P}_{\textsc{benign}}\bigr)$ to all participating clients. Given client $t$ with local dataset $\mathcal{D}$ and input data $x\in\mathcal{D}$, the client computes $q(x)$ and selects top-$N$ keys $\hat{\mathcal{K}}_{x}$ using Eq. 1 and update only the selected prompts using Eq. 2. If $x\equiv\mathcal{T}$, then we have the top-$N$ keys are exactly the the adversarial keys $\mathcal{K}_{\textsc{adv}}$, or $\hat{\mathcal{K}}_{x}\equiv\mathcal{K}_{\textsc{adv}}$ , which corresponds precisely to the membership condition.
@@ 3.2.3 Theoretical Analysis of PromptMIA
In this section, we theoretically analyze the performance of PromptMIA. Specifically, we study the True Positive Rate (TPR), the adversary’s success in identifying the target sample $\mathcal{T}$ when it is a member ($b=1$), and the False Positive Rate (FPR) which is the adversary’s error in identifying $\mathcal{T}$ when it is a non-member ($b=0$).
First, the PromptMIA attack is constructed to ensure perfect identification when the target sample is present. Theorem 1 establishes that the TPR of PromptMIA equals 1. The proof follows directly from the construction of adversarial keys in Algorithm 2, which ensures that $\max_{k_{a}\in\mathcal{K}_{\textsc{adv}}}\gamma(q(\mathcal{T}),k_{a})<\min_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\gamma(q(\mathcal{T}),k_{b})$. Thus, if the client processes $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected keys must be exactly the set $\mathcal{K}_{\textsc{adv}}$.
Theorem 1 (True Positive Rate).
@@
We now analyze the FPR, $\Pr[b^{\prime}=1\mid b=0]$, which is more complex. A false positive arises whenever a non-member query $q(x)$ selects all $N$ adversarial keys $\{k_{a_{j}}\}_{j=1}^{N}$ as its top-N choices than to any benign key $\{k_{b_{i}}\}_{i=1}^{M-N}$.
To bound this probability, we first model the distribution of non-member query $q(x)$ relative to the benign keys. The training objective in Eq. 2 includes a surrogate loss $\gamma(q(x),k_{i})$ that explicitly pulls a selected key $k_{i}$ to align with the query feature $q(x)$ that selected it. Thus, after several rounds of training, the benign keys $\mathcal{K}_{\textsc{benign}}$ will stabilize and function as the centroids for the query vectors $q(x)$ generated from the clients’ data (see Fig. 3 and Fig. 14). This observation forms our foundational assumptions:
@@
Next, we introduce the adversarial key into this setting. For a non-member query $q(x)$ drawn from cluster $i$, we define a cluster-flip event $E_{i}$ as the case when $q(x)$ selects all $N$ adversarial keys as its $N$ nearest centroids. Formally, $E_{i}$ is the intersection of $N\times(M-N)$ race events $A_{jl}$, where each $A_{jl}$ denotes that $q(x)$ is closer to an adversarial key $k_{a_{j}}$ than to a benign key $k_{b_{l}}$. A cluster-flip event thus arises only if the adversary wins all of these races simultaneously. The probability of the joint event $E_{i}$ can then be upper bounded by the probability of the single race with the lowest success probability, as formally stated in Lemma 1.
Lemma 1.
@@
With this lemma, we can bound the FPR. The FPR is the probability of the event $E_{FP}$ that that a single, random non-member query $q(x)$ results in a cluster-flip event. Using the Law of Total Probability, FPR can be expressed as a weighted sum of cluster-flip probabilities, where the weights correspond to the prior probabilities of clusters. Based on the bound on the probability of cluster-flip events for each cluster established in Lemma 1, we can derive a bound on the FPR which is the largest (worst-case) cluster-flip probability across all clusters, stated in Theorem 2.
Theorem 2 (False Positive Rate).
@@
where $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ is the set of $N$ adversarial keys, and $z_{ijl}$ is the $z$-score:
$z_{ijl}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$
Theorem 2 provides some insights on conditions under which the attack is most effective. First, the bound is tighter when the adversarial target $q(\mathcal{T})$ is highly distinctive. A distinctive target ensures that its corresponding adversarial keys are geometrically well separated from benign key clusters, resulting in small $\Phi\left(z_{ijl}\right)$. In addition, the FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys (i.e., minimizing $\sigma_{i}^{2}$). That will decreases the likelihood that any non-member query will randomly stray into the adversary’s region. Additionally, combining Theorems 1 and 2 yields the bound on the Advantage stated in Corollary 1. The detailed proofs of lemmas and theorems are provided in Appendix A.6.
Corollary 1 (Attack Advantage).
@@
3.3 Membership Inference Defenses against PromptMIA
Although the goal of this work is not to develop new defense mechanisms, we examine how standard approaches originally designed for gradient or output based MIAs interact with PromptMIA. A widely used defense is Differentially Private SGD (DPSGD) (Abadi et al., 2016; Duan et al., 2023), which clips per-sample gradients and injects noise before aggregation to provide $(\varepsilon,\delta)$-privacy guarantees. While DPSGD protects the content of gradients, in federated prompt tuning the top-$N$ prompt selection is independent of the gradient update procedure and therefore remains exposed, rendering DPSGD ineffective against PromptMIA. Therefore, we do not further evaluate DPSGD in our experiments. We consider Input Noise Perturbation, which adds calibrated noise directly to input pixels (Lecuyer et al., 2019). Similar to DPSGD, this approach incurs privacy-utility trade-off. Another line of defense is to use Anomaly Detection Methods (IsolationForest (Liu et al., 2008), LocalOutlierFactor (Breunig et al., 2000), OneClassSVM (Manevitz and Yousef, 2001), SGDOneClassSVM, and EllipticEnvelope (Rosseeuw, 1999) to filter out adversarial prompts. Although it would be interesting to evaluate deep learning–based anomaly detection algorithms (Do et al., 2025; Jiang et al., 2023) (e.g., autoencoders, VAEs, GAN-based models) against PromptMIA, we leave this for future work. Moreover, the number of prompts in the prompt pool is typically small, which raises doubts about the effectiveness of deep learning based methods in this setting. Details about these defenses are given in Appendix A.3. The effectiveness of anomaly detection and input noise perturbation against PromptMIA are reported in Sections 4.2 and 4.3, respectively. While our experimental results (Section 4.2) show that traditional outlier detection mechanisms fail to reliably identify adversarial keys, we take one step further and introduce a hyperparameter $\beta$ that controls the alignment between adversarial and benign keys. Increasing $\beta$ improves PromptMIA robustness against potential stronger anomaly detectors, albeit at the cost of reduced MIA accuracy (Appendix A.4).
4 Experimental Results
We evaluate PromptMIA on four datasets—CIFAR-10, CIFAR-100  (Krizhevsky et al., 2009), TinyImageNet  (Le and Yang, 2015), and a synthetic 4-dataset benchmark constructed by pooling MNIST-M (Lee et al., 2021), Fashion-MNIST  (Xiao et al., 2017), CINIC-10  (Darlow et al., 2018), and MMAFEDB  [^1] https://www.kaggle.com/datasets/yuulind/mmafedb-clean—and three different models: ViT-B32, ConViT and DeiT. Experiments follow four axes: (1) measuring attack effectiveness via advantage and success rate; (2) testing robustness of PromptMIA against classical anomaly detection methods (Isolation Forest, LOF, One-Class SVM, Elliptic Envelope); (3) analyzing the impact of input noise perturbation defenses; and (4) conducting ablations on key hyperparameters ($M$, NN, $\beta$, $\delta_{\min}$, $\Delta$). More information about experimental settings are given in Appx. A.5.
4.1 Advantage and Attack Success Rate Measurement
We evaluate the performance of PromptMIA against Federated Prompt Tuning using two metrics: Advantage (Eq. 3) and Attack Success Rate (Eq. 4). For all experiments, we set $\delta_{\text{min}}=0.02$ and $\Delta=0.05$. Unless otherwise specifically noted, we set $\beta=0$. Following (Wang et al., 2022), the global prompt pool size is fixed at $M=20$, and the prompt selection size at $N=4$. In the case of batched update, given batch $B=\{(x_{i},y_{i})\}_{i=1}^{\ell}$, for each sample the client computes the per-sample selected key set $\hat{\mathcal{K}}_{x_{i}}$ and corresponding per-sample loss $\mathcal{L}_{x_{i}}$ (defined in Eq. 2) and updates the batch-level set of chosen keys and prompts: $\hat{\mathcal{K}}_{B}=\hat{\mathcal{K}}_{B}\bigcup\hat{\mathcal{K}}_{x_{i}}$ and $\hat{\mathcal{P}}_{B}=\hat{\mathcal{P}}_{B}\bigcup\mathcal{P}_{x_{i}}$. Batch-wise loss $\mathcal{L}_{B}$ is calculated by accumulating $\mathcal{L}_{x_{i}}$. The client update the selected keys and prompts as $\hat{\mathcal{K}}_{B}\leftarrow\mathcal{K}_{B}-\mu\nabla_{\hat{\mathcal{K}}_{B}}\mathcal{L}_{B}$ and $\hat{\mathcal{P}}_{B}\leftarrow\hat{\mathcal{P}}_{B}-\mu\nabla_{\hat{\mathcal{P}}_{B}}\mathcal{L}_{B}$, where $\mu$ is the learning rate. After receiving the client’s updates, the server infers membership by checking whether all adversarial prompts are selected, i.e. $\mathbbm{1}_{\{\mathcal{T}\in\mathcal{D}\}}=1$ if $\mathcal{K}_{\textsc{ADV}}\subseteq\hat{\mathcal{K}}_{B}$, and 0 otherwise. Figure 4 shows average results of three models over four different datasets. Detailed results on individual models are given in Appendix A.7. PromptMIA consistently achieves near-perfect attack success rates across all models and datasets at small batch sizes, and maintains $90\%$ ASR and Advantage at larger batch sizes. In contrast, naive prompt injection collapses as the batch size increases, reaching Advantage of only $\approx 20\%$ against FourDataset when batch size =256=256.
@@ 4.1 Advantage and Attack Success Rate Measurement
4.2 Performance of outlier detection methods against PromptMIA
To assess the effectiveness of outlier detection against PromptMIA, we frame adversarial key detection as an unsupervised anomaly-detection task over the global prompt pool. For each input $x\in\mathcal{D}$, we flip $\tilde{b}\sim\mathrm{Bernoulli}(1/2)$: if $\tilde{b}=1$, $N$ adversarial keys are injected into the pool $\tilde{\mathcal{K}}$; otherwise, the pool remains unmodified (clean control). We then apply IsolationForest, LocalOutlierFactor, OneClassSVM, SGDOneClassSVM, and EllipticEnvelope to score keys, labeling injected ones as positives. Precision and Recall are computed using ground truth and averaged across all datasets and models (detailed results in App. A.8).

### v2

@@ 3.2.2 PromptMIA
To build up on and overcome the limitations of the naive attack, we impose two requirements on adversarial key generation. First, adversarial keys $\mathcal{K}_{\textsc{adv}}$ must achieve higher cosine similarity with the target query $q(\mathcal{T})$ than any benign key $\mathcal{K}_{\textsc{benign}}$. Second, adversarial keys must be sufficiently diverse from one another to not be trivially exposed. Formally, the server selects a subset $S\subseteq[M]$ of size $N$, and for each index $j\in S$, the original key $k_{j}$ is replaced with an adversarial key $k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}$ ( see Eq. 6). $k_{a_{m}}$ is constrained to lie within a cosine similarity interval with $q(\mathcal{T})$:
$\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)+\delta_{\min}\;\leq\;\kappa(q(T),k_{a_{m}})\;\leq$ (7)
$\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)+\delta_{\min}+\Delta,\quad\forall k_{a_{m}}\in\mathcal{K}_{\mathrm{adv}}.$
where $\delta_{\min}$ ensures that all adversarial keys have higher cosine similarity to $q(\mathcal{T})$ than benign keys, while $\Delta$ introduces controlled variability to prevent all adversarial keys from collapsing to the same vector. Algorithm 1 describes the procedure for generating a single adversarial key with a specified cosine similarity $s$ to the target query. Building on this primitive, Algorithm 2 constructs a set of $N$ adversarial keys $\mathcal{K}_{\textsc{ADV}}$ that satisfies Eq. 7. Details of these algorithms are given in Appendix A.2. PromptMIA allows the adversary to generate an adversarial key set $\mathcal{K}_{\textsc{adv}}$ that will always be selected and updated when $\mathcal{T}\in\mathcal{D}$, while not being easily detectable (see Fig. 2(c) and Section 4.2). The server generates and distributes the modified global prompt pool $(\tilde{\mathcal{K}},\tilde{\mathcal{P}})=\bigl(\mathcal{K}_{\textsc{adv}}\cup\mathcal{K}_{\textsc{benign}},\;\mathcal{P}_{\textsc{adv}}\cup\mathcal{P}_{\textsc{benign}}\bigr)$ to all participating clients. Given client $t$ with local dataset $\mathcal{D}$ and input data $x\in\mathcal{D}$, the client computes $q(x)$ and selects top-$N$ keys $\hat{\mathcal{K}}_{x}$ using Eq. 1 and update only the selected prompts using Eq. 2. If $x\equiv\mathcal{T}$, then we have the top-$N$ keys are exactly the adversarial keys $\mathcal{K}_{\textsc{adv}}$, or $\hat{\mathcal{K}}_{x}\equiv\mathcal{K}_{\textsc{adv}}$ , which corresponds precisely to the membership condition.
@@ 3.2.3 Theoretical Analysis of PromptMIA
In this section, we theoretically analyze the performance of PromptMIA. Specifically, we study the True Positive Rate (TPR), the adversary’s success in identifying the target sample $\mathcal{T}$ when it is a member ($b=1$), and the False Positive Rate (FPR) which is the adversary’s error in identifying $\mathcal{T}$ as a member when it is a non-member ($b=0$).
First, the PromptMIA attack is constructed to ensure perfect identification when the target sample is present. Theorem 3.1 establishes that the TPR of PromptMIA equals 1. The proof follows directly from the construction of adversarial keys in Algorithm 2, which ensures that $\min_{k_{a}\in\mathcal{K}_{\textsc{adv}}}\kappa(q(\mathcal{T}),k_{a})>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa(q(\mathcal{T}),k_{b})$. Thus, if the client possesses $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected keys must be exactly the set $\mathcal{K}_{\textsc{adv}}$.
Theorem 3.1 (True Positive Rate).
@@
We now analyze the FPR, $\Pr[b^{\prime}=1\mid b=0]$, which is more complex. A false positive arises whenever a non-member query $q(x)$ selects all $N$ adversarial keys $\{k_{a_{j}}\}_{j=1}^{N}$ as its top-N choices than to any benign key $\{k_{b_{i}}\}_{i=1}^{M-N}$. To bound this probability, we first model the distribution of non-member query $q(x)$ relative to the benign keys. The training objective in Eq. 2 includes a surrogate loss $\gamma(q(x),k_{i})$ that explicitly pulls a selected key $k_{i}$ to align with the query feature $q(x)$ that selected it. Thus, after several rounds of training, the benign keys $\mathcal{K}_{\textsc{benign}}$ will stabilize and function as the centroids for the query vectors $q(x)$ generated from the clients’ data (see Fig. 3 and Fig. 15). This observation forms our foundational assumptions:
@@
Next, we introduce the adversarial key into this setting. For a non-member query $q(x)$ drawn from cluster $i$, we define a cluster-flip event $E_{i}$ as the case when $q(x)$ selects all $N$ adversarial keys as its $N$ nearest centroids. Formally, $E_{i}$ is the intersection of $N\times(M-N)$ race events $A_{jl}$, where each $A_{jl}$ denotes that $q(x)$ is closer to an adversarial key $k_{a_{j}}$ than to a benign key $k_{b_{l}}$. A cluster-flip event thus arises only if the adversary wins all of these races simultaneously. The probability of the joint event $E_{i}$ can then be upper bounded by the probability of the single race with the lowest success probability, as formally stated in Lemma 3.2.
Lemma 3.2.
@@
where $\Phi(\cdot)$ is the CDF of the standard normal distribution.
With this lemma, we can bound the FPR. The FPR is the probability of the event $E_{FP}$ that that a single, random non-member query $q(x)$ results in a cluster-flip event. Using the Law of Total Probability, FPR can be expressed as a weighted sum of cluster-flip probabilities, where the weights correspond to the prior probabilities of clusters. Based on the bound on the probability of cluster-flip events for each cluster established in Lemma 3.2, we can derive a bound on the FPR which is the largest (worst-case) cluster-flip probability across all clusters, stated in Theorem 3.3.
Theorem 3.3 (False Positive Rate).
@@
$FPR\leq\max_{1\leq i\leq M-N}\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}\right)\right)$
where $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ is the set of $N$ adversarial keys, and $z_{ijl}$ is the $z$-score: $z_{ijl}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$
Theorem 3.3 provides some insights on conditions under which the attack is most effective. First, the bound is tighter when the adversarial target $q(\mathcal{T})$ is highly distinctive. A distinctive target ensures that its corresponding adversarial keys are geometrically well separated from benign key clusters, resulting in small $\Phi\left(z_{ijl}\right)$. In addition, the FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys (i.e., minimizing $\sigma_{i}^{2}$). That will decreases the likelihood that any non-member query will randomly stray into the adversary’s region. Additionally, combining Theorems 3.1 and 3.3 yields the bound on the Advantage stated in Corollary 3.4. The detailed proofs are provided in Appendix A.7.
Corollary 3.4 (Attack Advantage).
@@
3.3 MIA Defenses against PromptMIA
Although the goal of this work is not to develop new defense mechanisms, we examine how standard approaches originally designed for gradient or output based MIAs interact with PromptMIA. A widely used defense is Differentially Private SGD (DPSGD) (Abadi et al., 2016; Duan et al., 2023), which clips per-sample gradients and injects noise before aggregation to provide $(\varepsilon,\delta)$-privacy guarantees. While DPSGD protects the content of gradients, in federated prompt tuning the top-$N$ prompt selection is independent of the gradient update procedure and therefore remains exposed, rendering DPSGD ineffective against PromptMIA. Therefore, we do not further evaluate DPSGD in our experiments. We consider Input Noise Perturbation, which adds calibrated noise directly to input pixels (Lecuyer et al., 2019). Similar to DPSGD, this approach incurs privacy-utility trade-off. Another line of defense is to use Anomaly Detection Methods (IsolationForest (Liu et al., 2008), LocalOutlierFactor (Breunig et al., 2000), OneClassSVM (Manevitz and Yousef, 2001), and EllipticEnvelope (Rosseeuw, 1999) to filter out adversarial prompts. The effectiveness of anomaly detection and input noise perturbation against PromptMIA are reported in Sections 4.2 and 4.3, respectively. Although it would be interesting to evaluate deep learning-based anomaly detection algorithms (Do et al., 2025; Jiang et al., 2023) against PromptMIA, we leave this for future work. Moreover, the number of prompts in the prompt pool is typically small, which raises doubts about the effectiveness of deep learning based methods in this setting. Appendix A.3 presents a detailed discussion on theses defenses. We also analyzes the implications of other potential system-level defense strategies such as randomized key indices, secure aggregation (Bonawitz et al., 2017), and randomized prompt selection via prompt dropout and show that these defenses are either provably or empirically ineffective (Appx. A.4).
While our results (Section 4.2) show that traditional outlier detection fails to reliably identify adversarial keys, we further introduce a hyperparameter $\beta$ to control the alignment between adversarial and benign keys. Larger $\beta$ improves PromptMIA robustness against stronger anomaly detectors, at the cost of reduced MIA accuracy (Appendix A.5).
4 Experimental Results
We evaluate PromptMIA on four datasets: CIFAR-10, CIFAR-100  (Krizhevsky et al., 2009), TinyImageNet  (Le and Yang, 2015), and a synthetic 4-dataset benchmark constructed by pooling MNIST-M (Lee et al., 2021), Fashion-MNIST  (Xiao et al., 2017), CINIC-10  (Darlow et al., 2018), and MMAFEDB  [^1] https://www.kaggle.com/datasets/yuulind/mmafedb-clean and three different models: ViT-B32, ConViT and DeiT. Experiments follow four axes: (1) measuring attack effectiveness via advantage and success rate; (2) testing robustness of PromptMIA against classical anomaly detection methods; (3) analyzing the impact of input noise perturbation defenses; and (4) conducting ablations on key hyperparameters ($M$, NN, $\beta$, $\delta_{\min}$, $\Delta$). Details about experimental settings are given in Appx. A.6.
4.1 Advantage and Attack Success Rate Measurement
We evaluate the performance of PromptMIA against Federated Prompt Tuning using two metrics: Advantage (Eq. 3) and Attack Success Rate (Eq. 4). For all experiments, we set $\delta_{\text{min}}=0.02$ and $\Delta=0.05$. Unless otherwise specifically noted, we set $\beta=0$. Following (Wang et al., 2022), the global prompt pool size is fixed at $M=20$, and the prompt selection size at $N=4$. In the case of batched update, given batch $B=\{(x_{i},y_{i})\}_{i=1}^{\ell}$, for each sample the client computes the per-sample selected key set $\hat{\mathcal{K}}_{x_{i}}$ and corresponding per-sample loss $\mathcal{L}_{x_{i}}$ (defined in Eq. 2) and updates the batch-level set of chosen keys and prompts: $\hat{\mathcal{K}}_{B}=\hat{\mathcal{K}}_{B}\bigcup\hat{\mathcal{K}}_{x_{i}}$ and $\hat{\mathcal{P}}_{B}=\hat{\mathcal{P}}_{B}\bigcup\mathcal{P}_{x_{i}}$. Batch-wise loss $\mathcal{L}_{B}$ is calculated by accumulating $\mathcal{L}_{x_{i}}$. The client update the selected keys and prompts as $\hat{\mathcal{K}}_{B}\leftarrow\mathcal{K}_{B}-\mu\nabla_{\hat{\mathcal{K}}_{B}}\mathcal{L}_{B}$ and $\hat{\mathcal{P}}_{B}\leftarrow\hat{\mathcal{P}}_{B}-\mu\nabla_{\hat{\mathcal{P}}_{B}}\mathcal{L}_{B}$, where $\mu$ is the learning rate. After receiving the client’s updates, the server infers membership by checking whether all adversarial prompts are selected, i.e. $\mathbb{I}_{\{\mathcal{T}\in\mathcal{D}\}}=1$ if $\mathcal{K}_{\textsc{ADV}}\subseteq\hat{\mathcal{K}}_{B}$, and 0 otherwise. Figure 4 shows average results of three models over four different datasets. Detailed results on individual models are given in Appendix A.8. PromptMIA consistently achieves near-perfect attack success rates across all models and datasets at small batch sizes, and maintains $90\%$ ASR and Advantage at larger batch sizes. In contrast, naive prompt injection collapses as the batch size increases, reaching Advantage of only $\approx 20\%$ against FourDataset when batch size =256=256.
4.2 Performance of Outlier Detection Dethods
To evaluate outlier detection methods against PromptMIA, we frame adversarial key detection as an unsupervised anomaly-detection task over the global prompt pool. For each input $x\in\mathcal{D}$, we flip $\tilde{b}\sim\mathrm{Bernoulli}(1/2)$: if $\tilde{b}=1$, $N$ adversarial keys are injected into the pool $\tilde{\mathcal{K}}$; otherwise, the pool remains unmodified (clean control). We then apply IsolationForest, LocalOutlierFactor, OneClassSVM and EllipticEnvelope to score each key, where keys identified as outliers are considered adversarial. Results are averaged across all datasets and models (detailed results in App. A.9).
(a) CIFAR10
(b) CIFAR100
(c) TinyImagenet
(d) FourDataset